According to a Pew Research Center study, in May 2011, 33% of all American adults had a smart phone (one which the user can use to read email and surf the Internet). A communications professor at a university believes this percentage is higher among community college students. She selects 349 community college students at random and finds that 138 of them have a smart phone. In testing the hypotheses:

H0: p = 0.33 versus

Ha: p > 0.33,

she calculates the test statistic as z = 2.5990.

Then the p‑value = .

(Please round your answer to four decimal places.)
Answer #1

solution

z = 2.5990

P(Z>2.5990)=1 - P(Z<2.5990) = 1 -0.9953=0.0047(using z table)

p-value=0.0047